OpenStudy (anonymous):

Part 1: Find the perimeter of a rectangular object which has a length of squareroot 128 feet and a width of squareroot 200 feet. (3 points) Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point) show work

OpenStudy (amistre64):

the perimeter of a rectangle is adding up 4 sides 2 sides of sqrt128 each, plus 2 sides of sqrt200 each

OpenStudy (amistre64):

to simplify them, just determine their perfect square factors

OpenStudy (anonymous):

can you teach me how to determine their perfect square factors please

OpenStudy (fifciol):

let's take \[\sqrt{192}\] I can write it down in this way:\[\sqrt{192}=\sqrt{64*3}=\sqrt{64}*\sqrt3=8\sqrt3\] I cannot do that with 3 so I leave it in that form Can you do that with sqrt 128 and sqrt 200?

OpenStudy (anonymous):

sqrt 128= sqrt 64*2 = 8sqrt2? is that right :(

OpenStudy (fifciol):

yes

OpenStudy (fifciol):

what about sqrt200 ?

OpenStudy (anonymous):

sqrt200= sqrt100*2 = 10sqrt2? :(

OpenStudy (fifciol):

very good :)

OpenStudy (anonymous):

\[P=2\left( \sqrt{128} +\sqrt{200}\right)\] \[=2\left( \sqrt{64*2}+\sqrt{100*2} \right)\] \[=2\left( \sqrt{8^{2}*2}+\sqrt{10^{2}*2} \right)\] \[=2\left( 8\sqrt{2}+10\sqrt{2}\right)=2*18\sqrt{2}=36\sqrt{2}\] i think now it is clear.

OpenStudy (anonymous):

so basically you just add the outside numbers and keep the 2 inside the radical?

OpenStudy (anonymous):

and thats my perimeter?

OpenStudy (anonymous):

yes

OpenStudy (fifciol):

3 apples + 4 apples = 7 apples so 8sqrt2 + 10sqrt2=18 sqrt2

OpenStudy (anonymous):

you can add or subtract only if we have the same number under the radical sign.

OpenStudy (anonymous):

wait im confused is it 18sqrt2 or 36sqrt2

OpenStudy (fifciol):

the perimeter is 36sqrt2

OpenStudy (anonymous):

Thank you both so much i understand now (: